(b) -27(1+iV3) r= 12.56637 (C) (1 + i) o= 60 T=

Find the polar form, need help with theta for b and whole part of c
Answer #1

2 I 3 -207(1+i13) = -41(+iva --12.57(Cos I usne =-12-57 COSY+ 12.59)ising SOR=1-12.574 =12:57, O-5-610 R-112:57], 0= (1+iye =